# Smoothed Graph Contrastive Learning via Seamless Proximity Integration

### v1

@@ 2.3 Leveraging the advantages of graph geometry
In the context of contrastive learning on graphs, the positive matrix $\tilde{\Pi}_{pos}^{(i,j)}$ can be considered as a mapping from $\mathcal{G}^{(i)}$ to $\mathcal{G}^{(j)}$, with its rows and columns corresponding to nodes in $\mathcal{G}^{(i)}$ to $\mathcal{G}^{(j)}$, respectively. The goal of the smoothing approach is to extend this mapping to the neighbors of the paired nodes. In this specific context, since the rows of the positive matrix $\Pi_{pos}^{(i,j)}$ are associated with nodes in $\mathcal{G}^{(i)}$, the smoothing approach utilizes the geometry of graph view $\mathcal{G}^{(i)}$. Similarly, for the positive matrix $\Pi_{pos}^{(j,i)}$, the smoothing approach utilizes to the geometric properties of the graph view $\mathcal{G}^{(j)}$.
@@
Figure 2 illustrates the distinctions between positive and negative pairs in the conventional contrastive learning framework and our proposed smoothed contrastive approach. Notably, when considering a specific anchor node $v_{t}^{(i)}$ in $\mathcal{G}^{(i)}$ paired with $v_{k}^{(j)}$ in $\mathcal{G}^{(j)}$, the graph information from $\mathcal{G}^{(i)}$ is employed to generate the smoothed positive and negative pairs matrices $\tilde{\Pi}_{pos}^{(i,j)}$ and $\tilde{\Pi}_{neg}^{(i,j)}$.
@@ 3 Method: smoothed graph contrastive learning
3.1 Preliminaries
In the domain of unsupervised graph representation learning, we introduce an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ constitutes the node set $\{{v}_{1},{v}_{2},...,{v}_{N}\}$, and $\mathcal{E}$ denotes the edge set, formally captured as $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Within this contextual framework, we establish the definition of two pivotal matrices: the feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$, wherein each $\mathbf{x}_{i}\in\mathbb{R}^{F}$ represents the feature vector associated with a distinct node ${v}_{i}$; and the adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$, where the binary element $a_{i,j}$ equals $1$ if and only if an edge exists between nodes $v_{i}$ and $v_{j}$.
3.2 Framework
@@
The main objective of the proposed framework is unsupervised node classification, with a specific focus on effectively addressing medium and large-scale graphs. We introduce a novel framework, termed Smoothed Graph Contrastive Learning (SGCL), designed to construct node embeddings by seamlessly integrating the geometric structure of augmented graphs to facilitate a smooth alignment between positive and negative pairs.
Figure 3: Overview of the Proposed SGCL Model. The SGCL model utilizes a mini-batch generation approach to generate $|B|$ subgraphs (mini-batches). For each subgraph $\hat{\mathcal{G}}_{b}$, two distinct augmentation methods are employed to produce two different views $\hat{\mathcal{G}}_{b}^{(i)}$ and $\hat{\mathcal{G}}_{b}^{(j)}$. The GCN encoders $f_{{\theta}_{i}}$ and $f_{{\theta}_{j}}$ are then utilized to learn feature embeddings $\mathbf{H}_{b}^{(i)}$ and $\mathbf{H}_{b}^{(j)}$, respectively. Subsequently, two smoothed positive matrices $\tilde{\Pi}_{pos}^{(i,j)}$ and $\tilde{\Pi}_{pos}^{(j,i)}$ are computed by leveraging the geometric structures of the graphs. These matrices are used in the contrastive loss $\mathcal{J}_{SGCL}$ to encourage the smooth interaction of node representations between the two views. This loss function, scores the agreement between these representations for all batches, serving as training loss.
The comprehensive architecture of our framework is visually illustrated in Figure 3. In the following sections, we will outline the sequential processing steps of the proposed framework.
3.2.1 Subgraph generating
@@
To address the challenge of scalability and to accommodate the contrastive loss for large-scale graphs, we leverage the random-walk mini-batches generation approach, proposed by GraphSAINT (Zeng et al., 2020), to generate subgraphs from a given graph. More specifically, an entire graph $\mathcal{G}$ is partitioned into a set of $|B|$ mini-batches denoted as $\hat{\mathcal{G}}=\{\hat{\mathcal{G}}_{1},\ldots,\hat{\mathcal{G}}b,\ldots,\hat{\mathcal{G}}_{|B|}\}$, where each $\hat{\mathcal{G}}_{b}=(\hat{\mathcal{V}_{b}},\hat{\mathcal{E}_{b}})$ represents a sampled subgraph. It is essential to note that the construction of subgraphs varies depending on the specific sampling approach employed. Leveraging the insights gained from the variance analysis within GraphSAINT, it introduces a collection of lightweight and efficient mini-batch generation approaches, which are detailed in Appendix B.
3.2.2 Generating graph views via augmentation
Graph augmentation in GCL models involves two main approaches: (1) adjusting node features, which manipulates the initial node features, and (2) modifying graph structure, which includes alterations to the underlying structure of graphs. In the proposed framework, such graph augmentation strategies may pose challenges, especially in constructing positive and negative node pairs across multiple views. In the proposed framework, we employ a blend of edge-dropping and node feature masking strategies to generate two distinct graph views for every mini-batch $\hat{\mathcal{G}}_{b}$, denoted as $\hat{\mathcal{G}}_{b}^{(1)}$ and $\hat{\mathcal{G}}_{b}^{(2)}$.
@@ 3.2.2 Generating graph views via augmentation
More specifically, in each view $i$, we construct the augmented graph $\hat{\mathcal{G}}_{b}^{(i)}$ as $\hat{\mathcal{G}}_{b}^{(i)}=(\mathcal{T}_{x}^{(i)}(\mathbf{X}_{b}),\mathcal{T}_{A}^{(i)}(\mathbf{A}_{b}))$, where $\mathcal{T}_{x}(\mathbf{X})=\mathbf{X}\odot(1-M_{X})$ and $\mathcal{T}_{A}(\mathbf{A})=\mathbf{A}\odot(1-M_{A})+(1-\mathbf{A})\odot M_{A}$. Here, $M_{X}\sim\mathcal{N}(0,\Sigma)$ masks original values with Gaussian noise, and $M_{A}$ utilizes a Bernoulli distribution to randomly drop edges from the adjacency matrix.
3.2.3 Encoders
The encoder network with a parameter set $\theta$, denoted as $f_{\theta}$, processes an augmented graph as input, producing reduced-dimensional feature embeddings. Within this framework, we opt for the widely adopted Graph Convolution Network (GCN) (Kipf & Welling, 2017) as the foundational graph encoders. For each view $i$, we employ a dedicated graph encoder $\mathbf{H}=f_{{\theta}_{i}}(\mathbf{X},\mathbf{A}):\mathbb{R}^{N\times F}\times\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times F^{\prime}}$ that leverages adjacency and feature matrices as two congruent structural perspectives of GCN layers  [^1] For the sake of simplicity, we omit the view index in superscript and the batch index in subscript..
@@ 3.2.3 Encoders
The GCN operates across multiple layers, wherein the message-passing process is recurrently applied at each layer. The node representations are updated in a layer-wise manner: $\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$, where $\tilde{\mathbf{A}}$ denotes the symmetrically normalized adjacency matrix, calculated as $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ with diagonal matrix $\mathbf{I}\in\mathbb{R}^{N\times N}$, $\tilde{\mathbf{D}}_{ii}=\sum_{j}\tilde{\mathbf{A}}_{ij}\in\mathbb{R}^{N\times N}$ is the degree matrix, $\mathbf{W}^{(l)}\in\mathbb{R}^{F_{l}\times F_{l+1}}$ is the learned weight matrix for layer $l$, $\sigma$ is activation function, and $\mathbf{\mathbf{H}}^{(l)}\in\mathbb{R}^{N\times F_{l}}$ is the node representation in layer $l$.
3.2.4 Smoothed contrastive loss function
@@
To end-to-end training of the encoders and promote node representations, we introduce an innovative contrastive loss function. This loss function utilizes a smoothed positive pairs matrix $\tilde{\Pi}_{pos}^{(i,j)}$ to encourage the agreement between encoded embeddings of two nodes, namely, $v_{t}^{(i)}$ and $v_{p}^{(j)}$, in two different views with degree $\hat{\pi}_{pos}^{(i,j)}(t,p)$, while also distinguish their embeddings with a degree of $\hat{\pi}_{neg}^{(i,j)}(t,p)=1-\hat{\pi}_{pos}^{(i,j)}(t,p)$. The loss function is defined as follows:
@@
$\mathcal{L}_{SGCL}^{(i,j)}=$ $\parallel{\tilde{\Pi}_{pos}^{(i,j)}\odot(\mathbf{1}-\mathbf{C}^{(i,j)})}\parallel_{F}^{2}+$ (5)
$\lambda\parallel{\tilde{\Pi}_{neg}^{(i,j)}\odot\mathbf{C}^{(i,j)}}\parallel_{F}^{2}$
@@
where $\mathbf{1}$ is a matrix of the same size as $\mathbf{C}^{(i,j)}$ with all elements set to 1 and $\mathbf{C}^{(i,j)}$ is the Min-Max normalized cosine similarity matrix between the normalized embeddings $\mathbf{\hat{H}}^{(i)}$ and $\mathbf{\hat{H}}^{(j)}$ of identical networks:
$\mathbf{C}^{(i,j)}=Min-Max\left(\frac{\mathbf{\hat{H}}^{(i)}\mathbf{\hat{H}}^{{(j)}^{T}}}{\parallel\mathbf{\hat{H}}^{(i)}\parallel\parallel\mathbf{\hat{H}}^{(j)}\parallel}\right)$ (6)
@@
In the proposed contrastive loss function, the first term enforces the stability of the preservation in the embeddings of positive pairs by minimizing the discrepancy between 1 and each element of $\mathbf{C}^{(i,j)}$. This alignment is achieved with the values in the smoothed positive pairs matrix $\tilde{\Pi}_{pos}^{(i,j)}$, effectively equivalent to maximizing $\mathbf{C}^{(i,j)}$ for positive pairs. Conversely, the second term actively promotes a substantial diversity in the embeddings of negative pairs by minimizing each element of $\mathbf{C}^{(i,j)}$ concerning the values in the smoothed negative pairs matrix $\tilde{\Pi}_{neg}^{(i,j)}$.
At each training epoch, the smoothed positive pairs matrix $\tilde{\Pi}_{pos}^{(i,j)}$ is generated by applying various smoothing approaches outlined in Section 2.3. For instance, we use Taubin smoothing, which yields $\tilde{\Pi}_{pos}^{(i,j)}=\mathcal{S}_{T}(\Pi_{pos}^{(i,j)},\mathbf{L}^{(i)},\mu,\tau)$ and $\tilde{\Pi}_{pos}^{(j,i)}=\mathcal{S}_{T}(\Pi_{pos}^{(j,i)},\mathbf{L}^{(j)},\mu,\tau)$. The smoothed negative pairs matrix will be $\tilde{\Pi}_{neg}^{(i,j)}=1-\tilde{\Pi}_{pos}^{(i,j)}$. Ultimately, we learn the model parameters by considering all $|B|$ batches within the given graph concerning the overall innovated contrastive loss $\mathcal{J}_{SGCL}=\frac{1}{2|B|}\sum_{b=1}^{|B|}(\mathcal{L}_{SGCL}^{(i,j)}+\mathcal{L}_{SGCL}^{(j,i)})$.
In Equation 6, the hyperparameter $\lambda>0$ determines the trade-off between two terms during optimization. An ablation study evaluating the impact of different terms in the contrastive loss function, along with an exploration of the hyperparameter $\lambda$, is provided in Appendix E.2.

### v2

@@ 2.5 Smoothness promoting in positive and negative sets
In the context of contrastive learning on graphs, the positive matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ can be considered as a mapping from $\mathcal{G}^{(i)}$ to $\mathcal{G}^{(j)}$, with its rows and columns corresponding to nodes in $\mathcal{G}^{(i)}$ to $\mathcal{G}^{(j)}$, respectively. The goal of the smoothing approach is to extend this mapping to the neighbors of the paired nodes. In this specific context, since the columns of the positive matrix $\Pi_{\text{pos}}^{(i,j)}$ are associated with nodes in $\mathcal{G}^{(j)}$, the smoothing approach utilizes the geometry of graph view $\mathcal{G}^{(j)}$. Similarly, for the positive matrix $\Pi_{\text{pos}}^{(j,i)}$, the smoothing approach utilizes to the geometric properties of the graph view $\mathcal{G}^{(i)}$.
@@
Figure 2 illustrates the differences between positive and negative pairs in the conventional graph contrastive learning framework and our proposed smoothed contrastive approach. Notably, when considering a specific anchor node $v_{t}^{(i)}$ in $\mathcal{G}^{(i)}$ paired with $v_{k}^{(j)}$ in $\mathcal{G}^{(j)}$, the graph information from $\mathcal{G}^{(j)}$ is employed to generate the smoothed positive and negative pairs matrices $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ and $\tilde{\Pi}_{\text{neg}}^{(i,j)}$.
In the following, we analytically analyze the performance of smooth graph contrastive learning, by defining the following metrics.
Definition 2.3 (Dirichlet energy).
The Dirichlet energy of a signal $\mathbf{X}\in\mathbb{R}^{N\times F}$ on the vertices of a graph, defined as $E(\mathbf{X})=\mathbf{X}^{T}\mathbf{L}\mathbf{X}=\frac{1}{2}\sum_{i,j}a_{i,j}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}$, measures the smoothness of the signal $\mathbf{X}$ over the graph, where $\mathbf{L}=\mathbf{D}-\mathbf{W}$ is the graph Laplacian matrix.
A lower Dirichlet energy on a graph indicates that the signal $\mathbf{X}$ varies smoothly with minimal differences between adjacent nodes, aligning well with the graph structure as quantified by the graph Laplacian matrix [18].
Lemma 2.1 (Disparity).
For an encoder $f_{\theta}$, the disparity measure of learned features $\mathbf{X}\in\mathbb{R}^{N\times F}$ is defined by the distances of intra-class and inter-class Dirichlet energy as:
$D_{\text{disparity}}(f_{\theta})=\frac{1}{\mid E_{\text{intra}}\mid}\sum_{(i,j)\in E_{\text{intra}}}\Delta_{ij}-\frac{1}{\mid E_{\text{inter}}\mid}\sum_{(i,j)\in E_{\text{inter}}}\Delta_{ij}$
where $\Delta_{ij}=\frac{1}{2}a_{ij}\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|^{2}$, and $E_{\text{inter}}$ and $E_{\text{intra}}$ denote the sets of edges connecting nodes of different classes and within the same class, respectively. This measure captures the contrast in smoothness between intra-class and inter-class distances in the feature embedding.
A lower disparity measure indicates that the encoder produces node representations with greater similarity within the same class and increased distinction between different classes, reflecting more effective self-supervised learning.
Proposition 2.1.
For two graph encoders $f_{\theta}$ and $\tilde{f}_{\theta}$ learned using the conventional and smoothed graph contrastive frameworks, respectively, the disparity measure satisfies $D_{\text{disparity}}(\tilde{f}_{\theta})<D_{\text{disparity}}(f_{\theta})$.
This proposition indicates that geometry-aware graph contrastive losses enable the learned encoder to more effectively distinguish node representations.
We empirically validate this proposition by computing the disparity measure between two encoders, $f_{\theta}$ and $\tilde{f}_{\theta}$, used in the proposed SGCL and GRACE frameworks, respectively. Both encoders are applied to the same input graphs across a range of homophily rates. To ensure scale invariance in this comparison, we normalize the feature embeddings. This process removes the influence of the scales in the embeddings, allowing us to focus on the relative differences between embeddings rather than their absolute magnitudes. We use the graphs from [19], comprising 10 graphs with homophily rates $h$ varying from 0 to 0.9. Each graph contains 5000 nodes divided into two classes, sharing the same structure but differing in class labels. The results in Figure 3 indicate that the mean disparity of graph encoders used in all variants of SGCL is consistently lower than the conventional GCL approach reflecting a more effective self-supervised learning framework. Additionally, as $h$ increases, disparity measures decrease. This is because of the smoothing strategies that explore positive pairs in the proximity of each anchor node (and similarly for negative pairs). As the homophily rate increases, the number of false negatives inreases, and the role of SGCL in effectively contributing both positive and negative pairs to the contrastive loss becomes more prominent. Further analysis with real-world graphs can be found in Section E.1.
Figure 3: Comparison of mean disparity among graph encoders, illustrating that all SGCL variants consistently achieve lower values than the conventional GCL approach, which signifies a more effective self-supervised learning framework.
@@ 3 Method: smoothed graph contrastive learning
We introduce Smoothed Graph Contrastive Learning (SGCL), a novel framework that constructs node embeddings by seamlessly integrating the geometric structure of augmented graphs to ensure a smooth alignment between positive and negative pairs. The comprehensive architecture is illustrated in Figure 4. In the following sections, we outline the processing steps of the proposed framework.
Figure 4: Overview of the Proposed SGCL Model. The model first generates $|B|$ subgraphs and extracts two distinct views for each subgraph, denoted as $\hat{\mathcal{G}}_{b}^{(i)}$ and $\hat{\mathcal{G}}_{b}^{(j)}$. The GCN encoder is then employed to learn feature embeddings $\mathbf{H}_{b}^{(i)}$ and $\mathbf{H}_{b}^{(j)}$, respectively. Finally, the smoothed contrastive loss $\mathcal{L}_{\text{SGCL}}$ measures the agreement between these representations by utilizing $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ and $\tilde{\Pi}_{\text{neg}}^{(i,j)}$.
@@
Subgraph generating: We leverage the random-walk mini-batches generation approach [20] to generate subgraphs from a given graph. More specifically, an entire graph $\mathcal{G}$ is partitioned into a set of $|B|$ mini-batches denoted as $\hat{\mathcal{G}}=\{\hat{\mathcal{G}}_{1},\ldots,\hat{\mathcal{G}}b,\ldots,\hat{\mathcal{G}}_{|B|}\}$, where each $\hat{\mathcal{G}}_{b}=(\hat{\mathcal{V}_{b}},\hat{\mathcal{E}_{b}})$ represents a sampled subgraph. It is essential to note that the construction of subgraphs varies depending on the specific sampling approach employed. Leveraging the insights gained from the variance analysis within GraphSAINT [20], it introduces a collection of lightweight and efficient mini-batch generation approaches, further detailed in Appendix C.
@@
Generating graph views via augmentation: We employ a combination of edge-dropping and node feature masking strategies to generate two distinct graph views for every mini-batch $\hat{\mathcal{G}}_{b}$, denoted as $\hat{\mathcal{G}}_{b}^{(1)}$ and $\hat{\mathcal{G}}_{b}^{(2)}$. More specifically, in each view $i$, we construct the augmented graph $\hat{\mathcal{G}}_{b}^{(i)}$ as $\hat{\mathcal{G}}_{b}^{(i)}=(\mathcal{T}_{x}^{(i)}(\mathbf{X}_{b}),\mathcal{T}_{A}^{(i)}(\mathbf{A}_{b}))$, where $\mathcal{T}_{x}(\mathbf{X})=\mathbf{X}\odot(1-M_{X})$ and $\mathcal{T}_{A}(\mathbf{A})=\mathbf{A}\odot(1-M_{A})+(1-\mathbf{A})\odot M_{A}$. Here, $M_{X}\sim\mathcal{N}(0,\Sigma)$ masks original values with Gaussian noise, and $M_{A}$ utilizes a Bernoulli distribution to randomly drop edges from the adjacency matrix.
@@
Encoders: The encoder $f_{\theta}$ processes an augmented graph as input, producing reduced-dimensional feature embeddings. We choose the widely adopted Graph Convolutional Network (GCN) [2] as the graph encoder. For each view $i$, we employ a dedicated graph encoder $\mathbf{H}=f_{{\theta}_{i}}(\mathbf{X},\mathbf{A}):\mathbb{R}^{N\times F}\times\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times F^{\prime}}$ that leverages adjacency and feature matrices as two congruent structural perspectives of GCN layers  [^1] For the sake of simplicity, we omit the view index in superscript and the batch index in subscript.. The GCN operates across multiple layers, wherein the message-passing process is recurrently applied at each layer. The node representations are updated in a layer-wise manner: $\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$, where $\tilde{\mathbf{A}}$ denotes the symmetrically normalized adjacency matrix, calculated as $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ with diagonal matrix $\mathbf{I}\in\mathbb{R}^{N\times N}$, $\tilde{\mathbf{D}}_{ii}=\sum_{j}\tilde{\mathbf{A}}_{ij}\in\mathbb{R}^{N\times N}$ is the degree matrix, $\mathbf{W}^{(l)}\in\mathbb{R}^{F_{l}\times F_{l+1}}$ is the learned weight matrix for layer $l$, $\sigma$ is activation function, and $\mathbf{\mathbf{H}}^{(l)}\in\mathbb{R}^{N\times F_{l}}$ is the node representation in layer $l$.
@@
Smoothed contrastive loss function: To end-to-end training of the encoders and promote node representations, we introduce an innovative contrastive loss function. This loss function utilizes a smoothed positive pairs matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ to encourage the agreement between encoded embeddings of two nodes, namely, $v_{t}^{(i)}$ and $v_{p}^{(j)}$, in two different views with degree $\hat{\pi}_{\text{pos}}^{(i,j)}(t,p)$, while also distinguish their embeddings with a degree of $\hat{\pi}_{\text{neg}}^{(i,j)}(t,p)=1-\hat{\pi}_{\text{pos}}^{(i,j)}(t,p)$. The loss function is defined as:
@@
$\mathcal{L}_{\text{SGCL}}^{(i,j)}=\parallel{\tilde{\Pi}_{\text{pos}}^{(i,j)}\odot(\mathbf{1}-\mathbf{C}^{(i,j)})}\parallel_{F}^{2}+\lambda\parallel{\tilde{\Pi}_{\text{neg}}^{(i,j)}\odot\mathbf{C}^{(i,j)}}\parallel_{F}^{2}$ (5)
@@
where $\mathbf{1}$ is a matrix of the same size as $\mathbf{C}^{(i,j)}$ with all elements set to 1 and $\mathbf{C}^{(i,j)}$ is the normalized cosine similarity matrix between the normalized embeddings $\mathbf{\hat{H}}^{(i)}$ and $\mathbf{\hat{H}}^{(j)}$ of identical networks:
$\mathbf{C}^{(i,j)}=\frac{1}{2}\left(\frac{\mathbf{\hat{H}}^{(i)}\mathbf{\hat{H}}^{{(j)}^{T}}}{\parallel\mathbf{\hat{H}}^{(i)}\parallel\parallel\mathbf{\hat{H}}^{(j)}\parallel}+1\right)$ (6)
@@
In the proposed contrastive loss function, the first term enforces the stability of the preservation in the embeddings of positive pairs by minimizing the discrepancy between 1 and each element of $\mathbf{C}^{(i,j)}$. This alignment is achieved with the values in the smoothed positive pairs matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}$, effectively equivalent to maximizing $\mathbf{C}^{(i,j)}$ for positive pairs. Conversely, the second term actively promotes a substantial diversity in the embeddings of negative pairs by minimizing each element of $\mathbf{C}^{(i,j)}$ concerning the values in the smoothed negative pairs matrix $\tilde{\Pi}_{\text{neg}}^{(i,j)}$.
At each training epoch, the smoothed positive pairs matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}$ is computed by applying one of the smoothing approaches outlined in Section 2.4. For example, we apply Taubin smoothing, resulting in the smoothed positive pairs matrix $\tilde{\Pi}_{\text{pos}}^{(i,j)}=\mathcal{S}_{T}(\Pi_{\text{pos}}^{(i,j)},\mathbf{L}^{(i)};\mu,\tau)$, and the corresponding smoothed negative pairs matrix $\tilde{\Pi}_{\text{neg}}^{(i,j)}=1-\tilde{\Pi}_{\text{pos}}^{(i,j)}$. Ultimately, we learn the model parameters by considering all $|B|$ batches within the given graph concerning the overall innovated contrastive loss $\mathcal{J}_{\text{SGCL}}=\frac{1}{2|B|}\sum_{b=1}^{|B|}(\mathcal{L}_{\text{SGCL}}^{(i,j)}+\mathcal{L}_{\text{SGCL}}^{(j,i)})$.
In Equation (5), the hyperparameter $\lambda>0$ determines the trade-off between two terms during optimization. A comprehensive ablation study on the hyperparameters can be found in Appendix F.